The shafts are made of A36 steel and each one has a diameter of 80mm. Determine the angle of
rotation of gear D and the angle of rotation at point A.

0.6 m 150 mm 10 kNm D 200 mm 0.6 m E 150 mm 0.6 m 2 kN.m
Answer #1

( k obm A somm TokN E 2 ₂ kNm zoom + 0.6m ok o.fm Referring to the FOD of shatt coe, Em=0 (10 x 103 - (2x103 - Flo.) 20 Fro2The Polas moment of inertia of the shaft J = T d4 = T (2,0804 32 32 -6 J = 4,0 21 x 10 TABLAB GJ Take 6=75 Gle for A36 sted 6